# M2P2: A Multi-Modal Passive Perception Dataset for Off-Road Mobility in Extreme Low-Light Conditions

### v1

@@ VII Conclusions and Future Work
We present M2P2, a multi-modal passive perception dataset for off-road mobility facing a variety of lighting conditions and off-road terrain conditions. We open source our multi-modal sensor suite design, including thermal, event, and stereo-RGB cameras, two IMUs, GPS, and a LiDAR for ground truth. We present a streamlined multi-modal calibration procedure for infrared radiation, per-pixel brightness changes, RGB pixels, IMU readings, and point clouds. Our preliminary results show that off-road navigation with obstacle avoidance is possible through only passive perception in no-light conditions using end-to-end learning and classical planning.
As the first step toward fully passive perception for off-road mobility in extreme low-light conditions, this work opens up a new avenue of future research. Currently only thermal images and an off-the-shelf generic depth reconstruction model are used to generate 2D scans and costmaps for collision avoidance. However, our dataset can be used to fine-tune the model for better reconstruction results in off-road environments. Event images are yet to be leveraged, considering its low latency, high dynamic range, and low power consumption, potentially for high-speed, off-road maneuvers through darkness. Another interesting future direction is to enable other mobility tasks than simple obstacle avoidance, such as Visual Inertial Odometry [40, 41, 42], SLAM [43, 44, 45], and off-road kinodynamics modeling [46, 47, 48, 49, 50, 51], all with the purely passive modalities available from our multi-modal sensor suite and dataset.

### v2

@@ VII Conclusions and Future Work
This paper introduces M2P2, a novel multi-modal passive perception dataset specifically designed to address the challenges of off-road robot mobility in extreme low-light conditions. Unlike existing datasets, M2P2 uniquely combines thermal, event, and stereo RGB cameras, along with IMUs, GPS, and LiDAR for ground truth, providing a comprehensive representation of challenging off-road, low-light environments. We make the M2P2 dataset, along with our sensor suite design, publicly available to facilitate further research. We also present a robust multi-sensor calibration procedure, ensuring accurate data alignment across all modalities. Our initial experiments demonstrate that, even in complete darkness, off-road navigation, scene understanding, and vehicle state estimation are achievable using purely passive sensing.
While these initial experiments showcase the promise of individual modalities and limited fusion, the full realization of M2P2’s potential requires deeper exploration of advanced sensor fusion techniques and their application to a wider range of mobility tasks. As the first step toward fully passive perception for off-road mobility in extreme low-light conditions, this work opens up a new avenue of future research. Some of the areas that could benefit from M2P2 include Visual Inertial Odometry [41, 42, 43], SLAM [44, 45, 46], and off-road kinodynamics modeling [47, 48, 49, 50, 51, 52], all with the purely passive modalities available from our multi-modal sensor suite and dataset.